Solve Problems Using Linear Systems
10. White vinegar is a solution of acetic acid in water. There are two strengths of white vinegar-a 5% solution and a 10% solution. How many mililitres of each solution must be mixed to make 50 mL of a 9% vinegar solution?
Can you please help me with these problem? Thank you very much!!!

Answer by checkley77(12844)

.10x+.05(50-x)=.09*50
.10x+2.5-.05x=4.5
.05x=4.5-2.5
.05x=2
x=2/.05
x=40 ml of 10% vinegar is used.
50-40=10 ml of 5% vinegar is used.
Proof:
.10*40+.05*10=.09*50
4+.5=4.5
4.5=4.5